40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 9 01:29:26 2017, gautam, Update, LSC, DRMI Nosie Budget v3.0 7x
    Reply  Mon Oct 9 11:55:01 2017, Koji, Update, LSC, DRMI Nosie Budget v3.0 
       Reply  Mon Oct 9 22:18:34 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_dark.pdf
          Reply  Tue Oct 10 22:04:06 2017, rana, Update, LSC, AS55Q Dark Noise 
          Reply  Wed Oct 11 14:42:03 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
             Reply  Wed Oct 11 19:31:32 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
                Reply  Thu Oct 12 01:50:11 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_dark.png
                   Reply  Thu Oct 12 12:17:28 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_AS55_40m.pdf
                      Reply  Fri Oct 13 12:26:12 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_DemodNoises.pdf
                         Reply  Mon Oct 16 16:01:04 2017, gautam, Update, LSC, AS55Q Dark Noise DemodBoardwOscope.pdfDemodBoardwOscope_ASD.pdf
                            Reply  Tue Oct 17 19:31:53 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_TFs.pdfAS55_DemodNoises.pdf
    Reply  Tue Nov 7 17:45:05 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 SRCL_MICH_whitenoise_tf.pdfPRCL_MICH_whitenoise_tf.pdfMICH_aux.pdfC1NB_disp_40m_MICH_NB_2017-10-08.pdf
       Reply  Tue Nov 7 22:56:21 2017, gautam, Update, LSC, DRMI locking recovered DRMI_7Nov20178.png
          Reply  Wed Nov 8 00:28:16 2017, gautam, Update, LSC, Laser intensity coupling measurement attempt PSL_RIN.pdf
       Reply  Wed Nov 8 09:37:45 2017, rana, Update, LSC, DRMI Nosie Budget v3.1 
          Reply  Wed Nov 8 09:59:12 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 
             Reply  Wed Nov 15 01:37:07 2017, gautam, Update, LSC, DRMI low freq. nosie improved C1NB_disp_40m_MICH_NB_2017-11-15.pdfOL_A2L_couplings.pdf
                Reply  Thu Nov 16 00:53:26 2017, gautam, Update, LSC, DRMI noise sub-budgets C1NB_a2l_40m_MICH_NB_2017-11-15.pdfC1NB_aux_40m_MICH_NB_2017-11-15.pdf
Message ID: 13412     Entry time: Tue Nov 7 17:45:05 2017     In reply to: 13367     Reply to this: 13413   13415
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI Nosie Budget v3.1 

Some days ago, I had tried to measure the SRCL->MICH and PRCL->MICH cross couplings using broadband noise injected between 120-180 Hz, a frequency band chosen arbitrarily, in hindsight, I could have done a more broadband test. I've spent some time including the infrastructure to calculate "White-Noise TFs" in the noise budgeting code, where a transfer function is estimated by injecting a "broadband" excitation into a channel of interest, and looking at the resulting response in MICH. I figured this would be useful to estimate other couplings as well, e.g. laser intensity nosie, oscillator noise etc.

I estimate the transfer function of the coupling using the relation (MICH is the median ASD of the MICH error signal in the below expression, and similarly for PRCL)

|H_{cpl}| = \sqrt{\frac{|\mathrm{MICH}^{2}_{\mathrm{exc}} - \mathrm{MICH}^{2}_{\mathrm{quiet}}|}{|\mathrm{PRCL}^{2}_{\mathrm{exc}} - \mathrm{PRCL}^{2}_{\mathrm{quiet}}|}}

Attachments #1 and #2 show the spectra of the MICH, PRCL and SRCL signals during 'quiet' times and during the injection, while Attachment #3 shows the calculated coupling TFs using the above relation. These are significantly different (more than 10dB lower) than the numbers I reported in elog 13367, where the measurement was made using swept sine. As can be seen in the attached plots, the injected broadband excitation is visible above the nominal noise level, and I calculated the white noise TFs using ~5mins of data which should be plenty, so I'm not sure atm what to make of the answers from swept-sine and broadband injections being so different.

Attachment #4 shows the noise budget from the October 8 DRMI lock with the updated SRCL->MICH and PRCL->MICH couplings (assumed flat, extrapolated from Attachment #2 in the 120-180Hz band). If these updated coupling numbers are to be believed, then there is still some unexplained noise around 100Hz before we hit the PD dark noise. To be investigated. But if Attachment #4 is to be believed, it is not surprising that there isn't significant coherence between SRCL/PRCL and MICH around 100Hz.

Nov 8 1600: Updating NB to inculde estimated Oplev A2L.

Quote:
 

AUX coupling

This is the other find.

  • While chatting with Gabriele, he suggested measuring the SRCL->MICH and PRCL->MICH cross couplings.
  • I injected a signal in SRCL servo EXC channel, and adjusted amplitude till coherence in MICH_IN1 was good.
  • The actual TF measured was MICH_IN1 / SRCL_IN1 (so units of cts/ct).
  • My multiplying the in-lock PRCL and SRCL IN1 signals by these coupling coefficients (assumed flat in frequency for now, note that measurement was only made between 100Hz and 1kHz), I get the trace labelled "AUX coupling" in Attachment #1 (this is the quadrature sum for SRCL and PRCL couplings).
  • Also repeated for PRCL -> MICH coupling in the same way.
  • Measurements of these TFs and coherence are shown in Attachment #5 (again png screenshot because of DTT).
  • However, there is no significant coherence in MICH/SRCL or MICH/PRCL in this frequency range.

This seems to be limiting us from saturating the dark noise once the coil de-whitening is engaged. But lack of coherence means the mechanism is not re-injection of SRCL/PRCL sensing noise? Need to think about what this means / how we can mitigate it.


Attachment 1: SRCL_MICH_whitenoise_tf.pdf  247 kB  | Hide | Hide all
SRCL_MICH_whitenoise_tf.pdf
Attachment 2: PRCL_MICH_whitenoise_tf.pdf  247 kB  | Hide | Hide all
PRCL_MICH_whitenoise_tf.pdf
Attachment 3: MICH_aux.pdf  159 kB  | Hide | Hide all
MICH_aux.pdf
Attachment 4: C1NB_disp_40m_MICH_NB_2017-10-08.pdf  440 kB  Uploaded Wed Nov 8 16:01:23 2017  | Hide | Hide all
C1NB_disp_40m_MICH_NB_2017-10-08.pdf
ELOG V3.1.3-